The spacecraft went into safehold on the evening of Feb 5. As with previous safehold events, this caused the instruments to automatically turn off and point the spacecraft’s solar arrays towards the sun. It appeared very similar to the prior safehold events which AIM has experienced. The operations team immediately began recovery procedures and continued working the recovery through the weekend. By Feb 10th the spacecraft was in nominal operations and the instruments were in the process of being turned on.
SOFIE was operating nominally until the February 5th safehold event.
SOFIE meteoric smoke measurements were analyzed in an attempt to refine the identification of smoke composition. The results are inconclusive and point to errors in the ozone correction for band 3. The smoke paper is under revision.
The new motion correction algorithms have been implemented and tested. The data products, particularly H2O, show significant improvements in the events that previously had been seriously corrupted due to problems with the motion correction. V1.022 Level 1 signal processing is now underway. An additional Level 2 software improvement that will lead to improvements in dust analysis will also be added before full V1.022 processing begins in the next few days.
CIPS was operating nominally until the February 5th safehold event. It was turned on again Monday, 02/09/09.
The “daily daisies” for the current version of the retrieval algorithm, v3.20 (first produced in November 2008), were evaluated to determine their adequacy for posting to the web site. It was found that the daisies from all seasons had problems with too much background (clouds apparently present where none should be). Thus v3.20 daisies have been removed from the web site; older versions of the data are still available for seasons prior to SH 2008-2009. A new approach to plotting the daisies was developed and tested on a subset of the SH 2008-2009 data. This new approach successfully removed the excess background, so this season is currently being processed in its entirety so that new daisies can be posted. This is a stop-gap measure while the retrievals are evaluated to determine the underlying cause of the excess background.
A procedure to systematically evaluate all data levels is being designed. Parameters that need to be saved and analyses that need to be performed are being defined, as is a subset of the orbits which will be processed for initial evaluation. This evaluation procedure will enable us to more effectively find and diagnose processing and retrieval errors, with an initial goal of producing vetted, quantitative data products at all levels and for all seasons within several weeks.
CDE was turned on February 9th following the February 5th safehold event. CDE returned to normal operations and continues to take data.